# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v2

@@ 3.2 Evaluation metrics
We evaluate the method with the Classifier-2-Sample Test (C2ST) (Friedman, 2004; Lopez-Paz & Oquab, 2017), which allows evaluating the accuracy of posterior distributions in cases when the ground-truth posterior is known (Lueckmann et al., 2021). Here, a value close to 0.5 means that a classifier does not effectively distinguish the two distributions, and therefore, the posterior estimate is close to the ground-truth posterior. A value close to 1 means that the classifier can distinguish the distributions very well, indicating a poor posterior estimation. This framework is not applicable in most practical SBI settings, since it requires samples from the true posterior and is primarily used in toy examples (e.g., Section 4.1).
@@
The average Negative Log probability of the True Parameters (NLTP; $-\mathbb{E}[\log q(\boldsymbol{\theta_{o}}|\boldsymbol{x_{o}})]$ ) is a commonly reported metric in simulation-based inference for problems where the true posterior is unknown (Greenberg et al., 2019; Papamakarios & Murray, 2016; Durkan et al., 2020). In the limit of a large number of pairs $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$, the average over the log probability of each pair $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$ approaches the expected KL divergence between the estimated and the true posterior (up to a term that is independent of the estimated posterior), as shown in (Lueckmann et al., 2021).
@@ 4 Results
We evaluate the performance of our multifidelity approach to simulation-based inference on three tasks. We start with the Ornstein-Uhlenbeck process, for which the likelihood function is available in closed form, and follow with two challenging neuroscience problems with computationally expensive simulators and for which no likelihood is available: a multicompartmental neuron model, and a neural network model with synaptic plasticity.
4.1 Ornstein-Uhlenbeck process
Consider a drift-diffusion process of a particle starting at position $X(0)$ and drifting towards an equilibrium state. In this task, the Ornstein-Uhlenbeck (OU) process is the high-fidelity model of this phenomenon and has four parameters: the long-term mean $\mu$, volatility $\sigma$, rate of mean reversion $\gamma$, and $\mu_{\mathrm{offset}}$, which is the distance between the initial position $X(0)$ and $\mu$. As low-fidelity model, we chose i.i.d. samples from a Gaussian distribution, parametrized by $\mu_{\mathrm{L}}$ and $\sigma_{\mathrm{L}}$ (Fig. 2; Appendix B for details).
Figure 2: (A) High-fidelity model: Time-dependent Ornstein-Uhlenbeck process with four parameters. (B) Low-fidelity model: Time-independent i.i.d. Gaussian samples with two parameters.
The Ornstein-Uhlenbeck (OU) process has a tractable likelihood, which is given by
$p_{\text{exact}}(\boldsymbol{X}\mid\mu,\sigma,\gamma,\mu_{\mathrm{offset}})=p(X(0))\prod_{t=1}^{n}\frac{1}{\sqrt{\pi g}\sigma}\exp^{K_{t}},$ (3)
where K is defined as
$K_{t}=-\frac{1}{g\sigma^{2}}\left(\left(\mu-X_{t}\right)-\sqrt{1-\gamma g}\left(\mu-X_{t-1}\right)\right)^{2},$
$p(X(0))=\mathcal{N}(\mu+\mu_{\text{offset}},1)$ is the distribution over initial conditions, and $g=(1-\exp(-2\gamma\Delta T))/\gamma$ (details in Appendix B.1) (Kou et al., 2012).
[…]
@@ 4.1 Ornstein-Uhlenbeck process
To evaluate MF-NPE, we compared the estimated densities to the respective reference posterior, estimated from the exact likelihood with rejection sampling (Martino et al., 2018). We quantified the performance with the Classifier-2-Sample Test (C2ST) over 30 observations and 10 network initializations per observation.
First, we evaluated our multifidelity approach in a setting where the low- and high-fidelity models have the same number of parameters (Fig. 3). Second, we gradually increased the number of parameters of the high-fidelity model by adding the mean reversion $\gamma$ and $\mu_{\mathrm{offset}}$, resulting in a setup with three and four parameters for the high-fidelity model versus two for the low-fidelity model. We designate the three experiments OU2, OU3 and OU4, respectively.
@@
Across experiments, we observed a consistent increase in performance with MF-NPE compared to NPE, especially in low simulation budgets from the high-fidelity model (50-1000 simulations) (Fig. 3). In addition, we found that having a higher number of low-fidelity samples improved performance, reinforcing that low-fidelity simulations were indeed advantageous for pre-training the neural density estimator for the downstream task. Note that in this task, we did not observe a substantial increase in MF-NPE performance between the setting with $10^{4}$ and $10^{5}$ low-fidelity samples, suggesting that we might have reached an upper bound regarding pre-training efficacy.
We expected that adding parameters to the high-fidelity model that are not present in the low-fidelity model, increased the complexity of the inference problem for MF-NPE, and indeed observed a decrease in MF-NPE performance, although MF-NPE still performed better than NPE.
@@
As described in Section 3, we enhanced the sequential algorithm TSNPE (Deistler et al., 2022) with a first round of MF-NPE, and designated this approach as MF-TSNPE. We found that MF-TSNPE (with 5 rounds; more details in Appendix F.1) performs better than the amortized MF-NPE, especially in regimes with low budget of high-fidelity simulations, further confirming previous findings that sequential methods have substantially higher performance in non-amortized scenarios (Lueckmann et al., 2021).
Finally, we found that A-MF-TSNPE shows minor improvements compared to MF-TSNPE. Such an approach is substantially more costly to run (given the ensemble of density estimators) and has a larger number of hyperparameters to tune. We cannot exclude the hypothesis that larger performance gains could be obtained from such an approach by a more extensive hyperparameter search.
Overall, the results of the Ornstein-Uhlenbeck task suggest that MF-NPE and MF-TSNPE have the potential to lead to substantial performance gains compared to NPE.
@@ 4.2 Multicompartmental neuron model
The voltage response of a morphologically-detailed neuron to an input current is typically modeled with a multicompartment model wherein the voltage dynamics of each compartment are based on the Hodgkin-Huxley equations (Hodgkin & Huxley, 1952; Hay et al., 2011; Druckmann et al., 2007). The higher the number of compartments of the model, the more accurate the model is, but the higher the simulation cost.
@@
To simulate the neuron models, we used Jaxley, a Python toolbox that makes use of GPU acceleration for efficiently simulating multicompartment single neurons with biophysical detail (Deistler et al., 2024). In this setting, the simulation time for the high-fidelity model is approximately 7 times higher than that of the low-fidelity model.
We characterized the neural response with four summary statistics that have been commonly used when fitting biophysical models of single neurons to empirical data: number of spikes, mean resting potential, standard deviation of the resting potential, and the voltage mean (Gonçalves et al., 2020; Gao et al., 2023).
The performances of MF-NPE and NPE were evaluated with NLTP on $10^{3}$ pairs of $\theta_{o}$ and respective simulation outputs $x_{o}$ (see Section 3.2), averaged over 10 different random network initialization and fixed simulation data set.
In line with the results in the OU process task, MF-NPE showed much better performance than NPE in terms of NLTP, in particular with higher low-fidelity simulation budgets (Fig. 4B,C). Furthermore, posterior predictives of MF-NPE closely matched the empirical data, in contrast with NPE, even when NPE was trained on a higher number of high-fidelity simulations (Fig. 4D; Appendix C). Finally, we ran simulation-based calibration tests and found that both MF-NPE and NPE estimates are relatively well calibrated (Fig. 4E) (Talts et al., 2020).
Figure 4: (A) A thick-tufted layer 5 pyramidal cell from the neocortex, containing 1 or 8 compartments per branch, respectively for the low- and high-fidelity models. (B) NLTP evaluated on 1000 pairs of observations and respective true parameters. Different colors of MF-NPE correspond to different numbers of low-fidelity samples used to pre-train the model (respectively $10^{3},10^{4},10^{5}$; same naming convention as in Fig. 3). (C) Posterior distributions across 4 methods, trained on $10^{3}$ high-fidelity samples (dotted box in panel B). (D) Posterior predictives, where NPE and MF-NPE have been trained on $10^{3}$ high-fidelity samples. (E) Simulation-based calibration for NPE and MF-NPE trained on $10^{3}$ samples.
@@ 4.3 Recurrent Spiking Network
Finally, we applied MF-NPE to a particularly challenging and timely problem in neuroscience: the inference of synaptic plasticity rules that endow large spiking neural networks with dynamics reminiscent of experimental data. This problem has been recently tackled with an SBI method (filter simulation-based inference, fSBI) that progressively narrows down the search space of parameters given different sets of summary statistics (Confavreux et al., 2023). fSBI was successful in obtaining manifolds of plasticity rules that ensure plausible network activity, but the compute requirements were reported to be very large. Here, we aim to test whether this problem can be efficiently tackled with MF-NPE.
@@
The high-fidelity simulator consisted of a recurrent network of 4096 excitatory ($E$) and 1024 inhibitory ($I$) leaky integrate-and-fire neurons connected with conductance-based synapses (Fig. 5A, B). Each synapse type in this network ($E$-to-$E$, $E$-to-$I$, $I$-to-$E$ and $I$-to-$I$) was plastic with an unsupervised local learning rule. For each synapse type, six parameters governed how the recent pre- and post-synaptic activity were used to update the synapse, for a total of 24 free parameters across all 4 synapse types, following previous work (Confavreux et al., 2023). The networks were simulated using Auryn, a C++ simulator (Zenke & Gerstner, 2014) (additional details in Appendix D).
Figure 5: (A) Example raster plots of the neural activities in a recurrent plastic network over 1s time window. (B) Schematics of the low- and high-fidelity models of a spiking network. (C) Performance of NPE and MF-NPE evaluated on 1000 true observations with NLTP, and averaged over 10 network initializations, with 95% confidence intervals.
@@
Mean-field theory can be applied to the dynamical system above to obtain the steady-state activities of the excitatory and inhibitory populations as a function of the parameters of the plasticity rules embedded in the network. Though such analysis is widely performed in the field (Vogels et al., 2011; Confavreux et al., 2023; Gerstner & Kistler, 2002; Gerstner et al., 2014), it has never been used as a low-fidelity model to help with the inference of the high-fidelity model parameters (additional details in Appendix D). Since there are no dynamics to simulate with the mean-field model, the simulation was almost instantaneous, while the high-fidelity model took approximately 5 minutes to generate a 2-minute long simulation on a single CPU.
@@
Summary statistics for both the low- and high-fidelity models were two-dimensional: the average firing rates of the excitatory and inhibitory neurons at steady state (after 2 minutes of simulation in the high-fidelity model). Plastic networks were considered plausible if the measured firing rates were between 1 and 50Hz (Dayan & Abbott, 2001; Confavreux et al., 2023).
Note that in our task, the low-fidelity model focuses solely on the $E$-to-$E$ and $E$-to-$I$ rules from the high-fidelity model, thereby having 12 out of the 24 parameters of the high-fidelity model. This setup allows us to demonstrate the performance of MF-NPE on problems with different parameter spaces, highlighting MF-NPE’s flexibility and advantages.
Table 1: Comparison between methods. MF-NPE4 is pretrained on $10^{4}$ low-fidelity samples, and MF-NPE5 on $10^{5}$ samples.
Method ##\## simulations rates 1–50 Hz
NPE HF: $10^{3}$ 66.4%
MF-NPE4 HF: $10^{3}$; LF: $10^{4}$ 94.2%
MF-NPE5 HF: $10^{3}$; LF: $10^{5}$ 95.6%
@@
We found that MF-NPE has better performance than NPE in terms of NLTP (Fig. 5C). Furthermore, MF-NPE leads to an increase of almost $30\%$ in the proportion of posterior samples within the target firing rate bounds (Table 1), reinforcing that MF-NPE is a practical and effective method for simulation-based inference of costly real-world simulators.

### v3

@@ 3.1.4 Evaluation metrics
We evaluate the accuracy of posterior distributions in cases where the ground-truth posterior is known with the Classifier-2-Sample Test (C2ST) and the Maximum Mean Discrepancy (MMD)(Friedman, 2004; Lopez-Paz & Oquab, 2017; Gretton et al., 2012; Lueckmann et al., 2021; Peyré & Cuturi, 2017). C2ST is commonly used in SBI, as it is easy to apply and interpret: a value close to 0.5 means that a classifier cannot effectively distinguish the two distributions, implying the posterior estimate is close to the ground-truth posterior. A value close to 1 means that the classifier can distinguish the distributions very well, indicating a poor posterior estimation. C2ST is rarely applicable in practical SBI settings, since it requires samples from the true posterior (e.g., Sec. 4.1).
@@
The average Negative Log probability of the True Parameters (NLTP; $\mathbb{E}[\log q(\boldsymbol{\theta_{o}}|\boldsymbol{x_{o}})]$ ) has been extensively used in the SBI literature for problems where the true posterior is unknown (Greenberg et al., 2019; Papamakarios & Murray, 2016; Durkan et al., 2020; Hermans et al., 2020). In the limit of a large number of pairs $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$, the average over the log probability of each pair $(\boldsymbol{\theta_{o}},\boldsymbol{x_{o}})$ approaches the expected KL divergence between the estimated and the true posterior (up to a term that is independent of the estimated posterior), as shown in (Lueckmann et al., 2021). In addition, we report the Normalized Root Mean Square Error (NRMSE), which quantifies the deviation of posterior samples from the true parameters on a scale-invariant axis. NRMSE values closer to 0 indicate better predictive performance.
@@ 4 Results
We evaluate the performance of our multifidelity approach to NPE and TSNPE on six tasks involving various types of observations (e.g., time series, images, neural spiking). We start with four benchmarking tasks, followed by two challenging neuroscience problems with computationally expensive simulators and for which no likelihood is available: a multicompartmental neuron model and a neural network model with synaptic plasticity. We also provide a comparison to MF-ABC (Sec. F.1.1, E.3).
4.1 Benchmarking tasks
We first evaluated MF-(TS)NPE on four benchmarking tasks: SIR, SLCP, OUprocess, and Gaussian Blob. SIR and SLCP are established SBI benchmarks (Lueckmann et al., 2021), OUprocess is a new multifidelity task with tractable likelihood (Kou et al., 2012), and Gaussian Blob is a high-dimensional image task (Lueckmann et al., 2019) (details in Appendix E). These tasks were chosen to systematically investigate various task properties that might impact the performance of transfer learning in a multifidelity setting: differing parameter dimensionality between the low- and high-fidelity models, partly observed dynamics, differing simulator types between the low- and high-fidelity models, and high-dimensional observations. Furthermore, these multifidelity tasks are not trivial in the sense that the low and high-fidelity simulators lead to different posteriors (Appendix J). Note that we do not evaluate the total cost of low- and high-fidelity simulations in these tasks, but defer this analysis to the two complex neuroscience tasks (Appendix K).
@@ 4.1 Benchmarking tasks
To evaluate MF-NPE, we compared the estimated densities to the respective reference posterior, estimated from the exact likelihood with Rejection Sampling (Martino et al., 2018) (OU process; closed-form of the likelihood in Sec. E.1), and using Sampling and Importance Resampling (RUBIN, 1988) to obtain a set of 10k proposal samples (SLCP, SIR), similar to Lueckmann et al. (2021). We quantified the performance with C2ST and MMD over 10 observations (30 observations for the OU process) and 10 network initializations per observation. GaussianBlob uses a CNN embedding and was evaluated with NRMSE and NLTP since no closed-form likelihood is available (Fig. 10).
Figure 2: (A)-MF-(TS)NPE outperforms NPE and TSNPE in simulation-efficiency. C2ST and MMD averaged over 10 network initializations with means and $95\%$ confidence intervals. MF-NPE4 and MF-NPE5 are pretrained on $10^{4}$ and $10^{5}$ low-fidelity simulations, respectively. Results for the GaussianBlob task in Fig. 10; variations on the OU task and comparisons to MF-ABC in Fig. 8.
@@
Across experiments, we observed a consistent performance increase with MF-NPE compared to NPE, and (A)-MF-TSNPE compared to TSNPE, especially in low simulation budgets from the high-fidelity model (50-$10^{3}$ simulations) (Fig. 2; Gaussian Blob in Fig. 10). In addition, we found that having a higher number of low-fidelity samples improved performance, reinforcing that low-fidelity simulations were indeed advantageous for pre-training the neural density estimator for the downstream task. Note that for the OU and SLCP tasks, we did not observe a substantial increase in MF-NPE performance between the settings with $10^{4}$ and $10^{5}$ low-fidelity samples, suggesting an upper bound regarding pre-training efficacy. We also compared MF-NPE with MF-ABC, an ABC-based method for multifidelity SBI (Prescott & Baker, 2020), and observed that MF-NPE has a substantially higher performance (Appendix F.1.1). This is consistent with previous findings indicating the superior performance of NPE with respect to rejection ABC and SMC-ABC, where it is not uncommon to require orders of magnitude more simulations to obtain reliable posterior approximations (Lueckmann et al., 2021; Frazier et al., 2024). However, a more extensive hyperparameter search could potentially lead to substantial improvements in MF-ABC performance.
@@
As described in Sec. 3, we enhanced the sequential algorithm TSNPE (Deistler et al., 2022) with a first round of MF-NPE, and designated this approach as MF-TSNPE. We found that MF-TSNPE (details in Appendix M.1) performs better than TSNPE, especially in regimes with a low budget of high-fidelity simulations. Compared to MF-TSNPE, A-MF-TSNPE improved inference in the OU process, but did not show significant improvements in the SLCP and SIR tasks.
Finally, we assessed the contribution of transfer learning to the overall performance in a setting where the low- and high-fidelity models have a different number of parameters, in the context of the OUprocess task (Appendix E.3). We expected that adding parameters to the high-fidelity model that are absent in the low-fidelity model would increase the inference complexity for MF-NPE, and indeed observed a performance decrease in MF-NPE, although MF-NPE still performed better than NPE and MF-ABC (see Appendix E.3). Overall, the results suggest that MF-NPE and MF-TSNPE can yield substantial performance gains compared to NPE, TSNPE, and MF-ABC.
@@ 4.2 Multicompartmental neuron model
The voltage response of a morphologically-detailed neuron to an input current is typically modeled with a multicompartment model wherein the voltage dynamics of each compartment are based on the Hodgkin-Huxley equations (Hodgkin & Huxley, 1952). The higher the number of compartments of the model, the more accurate the model is, but the higher the simulation cost.
@@
To simulate the neuron models, we used Jaxley, a Python toolbox for efficiently simulating multicompartment single neurons with biophysical detail (Deistler et al., 2024). In this setting, the simulation time for the high-fidelity model is approximately 4 times higher than that of the low-fidelity model. We characterized the neural response with four summary statistics that have been commonly used when fitting biophysical models of single neurons to empirical data: spike count, mean resting potential, standard deviation of the resting potential, and voltage mean (Gonçalves et al., 2020; Gao et al., 2023). Performances were evaluated with NLTP and NRMSE on $10^{3}$ pairs of $\theta_{o}$ and respective simulation outputs $x_{o}$, averaged over 10 random network initializations (Sec. 3.1.4).
MF-(TS)NPE showed higher performance than NPE, in particular with larger low-fidelity simulation budgets (Fig. 3B; Fig. G.1), despite the right-skewed posterior distribution of the low-fidelity model (Fig. 19). Furthermore, MF-NPE posterior predictives closely matched the empirical data, in contrast with NPE, even when NPE was trained on a higher number of high-fidelity simulations (Appendix G). In addition, MF-(TS)NPE achieved comparable performance with a total computational cost 4-5 times lower than standard NPE (Appendix K). Finally, simulation-based calibration tests suggest that both MF-NPE and NPE estimates were relatively well calibrated (Fig. 3C) (Talts et al., 2020).
(A)-MF-TSNPE pre-trained on $10^{4}$ low-fidelity samples outperforms MF-NPE trained on $10^{5}$ samples. However, A-MF-TSNPE performance comes at the cost of training time due to the use of an ensemble of density estimators (Appendix K). This additional training burden is only justified when the simulation cost is substantially higher than the training cost.
Figure 3: MF-NPE improves inference on neuroscience tasks. (A) Thick-tufted layer 5 pyramidal cell from the neocortex. (B) Performance evaluation with NLTP (same naming convention as in Fig. 2). Amortized methods are averaged over 10 network initializations; non-amortized trained once per 100 observations. Similar results were obtained with NRMSE (Appendix G.1). MF-NPE, and especially its sequential variants, are orders of magnitude more simulation-efficient than NPE. (C) Simulation-based calibration for NPE and MF-NPE trained on $10^{3}$ high-fidelity samples. Posterior samples and predictives are in Appendix G. (D) Schematic of the low and high-fidelity models of a spiking network. (E) Performance of NPE and MF-NPE evaluated on 10000 true observations with NLTP: averages over 10 network initializations, and 95% confidence intervals. (F) Proportion of posterior samples within the target firing rate bounds. MF-NPE produces a higher fraction of parameter sets within the bounds than NPE.
@@ 4.3 Recurrent spiking network
Finally, we applied MF-NPE to a challenging and timely problem in neuroscience: the inference of synaptic plasticity rules that endow large spiking neural networks with dynamics reminiscent of experimental data. This problem has been recently tackled with an SBI method (filter simulation-based inference, fSBI) that progressively narrows down the search space of parameters given different sets of summary statistics (Confavreux et al., 2023). fSBI was successful in obtaining manifolds of plasticity rules that ensure plausible network activity, but the compute requirements were reported to be very large. Here, we aim to test whether this problem can be efficiently tackled with MF-NPE.
@@
The high-fidelity simulator consisted of a recurrent network of 4096 excitatory ($E$) and 1024 inhibitory ($I$) leaky integrate-and-fire neurons connected with conductance-based synapses (Fig. 3D). Each synapse type in this network ($E$-to-$E$, $E$-to-II, II-to-$E$, II-to-II) was plastic with an unsupervised local learning rule. For each synapse type, 6 parameters governed how the recent pre- and post-synaptic activity were used to update the synapse, for a total of 24 free parameters across all 4 synapse types (Confavreux et al., 2023). The networks were simulated using Auryn, a C++ simulator (Zenke & Gerstner, 2014) (details in Appendix H).
@@
Mean-field theory can be applied to the dynamical system above to obtain the steady-state activities of the excitatory and inhibitory populations as a function of the parameters of the plasticity rules embedded in the network. Though such analysis is widely performed in the field (Vogels et al., 2011; Confavreux et al., 2023; Gerstner et al., 2014), it has never been used as a low-fidelity model to help with the inference of the high-fidelity model parameters. Since there are no dynamics to simulate with the mean-field model, the simulation was almost instantaneous, while the high-fidelity model took approximately 5 minutes to generate a single 2-minute long simulation on a single CPU.
@@
Summary statistics of the low- and high-fidelity models were the average firing rates of the excitatory and inhibitory neurons at steady state (after 2 minutes of simulation in the high-fidelity model). Plastic networks were considered plausible if the firing rates were between 1 and 50Hz (Dayan & Abbott, 2001; Confavreux et al., 2023).
@@
In this task, the low-fidelity model focuses solely on the $E$-to-$E$ and $E$-to-II rules from the high-fidelity model, thereby having 12 out of the 24 parameters of the high-fidelity model. This setup allows us to demonstrate the performance of MF-NPE on problems with different parameter spaces, highlighting MF-NPE’s flexibility and advantages. We found that MF-NPE has better performance than NPE in terms of NLTP (Fig. 3E). Furthermore, MF-NPE leads to an increase of almost $30\%$ in the proportion of posterior samples within the target firing rate bounds (Fig. 3F), reinforcing that MF-NPE is a practical and effective method for SBI of costly real-world simulators.
4.4 When does pre-training help?
In previous sections, we demonstrated that MF-NPE can significantly reduce the number of high-fidelity simulations required to accurately approximate the high-fidelity posterior by leveraging pre-training on low-fidelity simulations. This naturally leads to several key questions: Which characteristics of low-fidelity simulators enable effective transfer learning? Under what conditions can pre-training reliably enhance simulation efficiency?
Providing theoretical guarantees for these questions necessitates a formal characterization of convergence rates in NPE with transfer learning. Although recent works have begun addressing these challenges in NPE (Frazier et al., 2024), current theoretical frameworks of transfer learning (Tahir et al., 2024; Yun et al., 2020; Tripuraneni et al., 2020; Lampinen & Ganguli, 2018), rely on simplifying assumptions (e.g., linear networks) that do not fully capture the complexities of MF-NPE. Given this limitation, we instead empirically explored the conditions in which low-fidelity pre-training facilitates effective transfer learning. To do this, we evaluate MF-NPE where the low- and high-fidelity simulators are related by systematic perturbations (e.g., additive noise or a shift in representation) (Fig. 4). Through these experiments, we identified mutual information and representational similarity as two key dimensions in predicting the effectiveness of pre-training (details in Appendix B).